ganeshie8 (ganeshie8):

show that \(\sin(x)+\sin(x-120^{\circ})+\sin(x-240^{\circ})=0\)

ganeshie8 (ganeshie8):

my roomate and i are arguing about electric motors and im looking for some geometric explanation for this..

OpenStudy (empty):

what's true for x=0 is always true no matter how you rotate it

OpenStudy (empty):

\[e^{i x}(e^{i 2 \pi / 3} +e^{i 2 \pi 2/ 3} +e^{i 2 \pi 3/ 3}) =0\] the real part of that

OpenStudy (empty):

imaginary part I mean, whatever, both are zero haha

OpenStudy (michele_laino):

it is a three phase system

OpenStudy (rhr12):

x + 120 = t sin(t - 120) + sin(t) + sin(t + 120) = 0 sin(t)cos(120) - sin(120)cos(t) + sin(t) + sin(t)cos(120) + sin(120)cos(t) = 0 2 * cos(120) * sin(t) + sin(t) = 0 sin(t) * (2 * cos(120) + 1) = 0 sin(t) * (2 * (-1/2) + 1) = 0 sin(t) * (-1 + 1) = 0 sin(t) * 0 = 0 So, sin(t) can equal pretty much anything t = x + 120 sin(x + 120) can be anything x + 120 can be anything x can be anything There are no wrong solutions for this one.

OpenStudy (rhr12):

That's all i got.

ganeshie8 (ganeshie8):

Yes he is saying delta/star and some bs which i have no idea about... but it makes sense to think in terms of roots of unity as Empty wa suggesting.. (if im interepreting correctly..)

OpenStudy (michele_laino):